Say the demand equation is given as: Qd = -Price +2 The supply equation is Qs = 2Price. The equilibrium price is

a. 3/2

b. 2

c. 3

d. 2/3
Answer #1

Equllibrium price will be known

Qd = Qs

2-P = 2P

2 = 3P

P = 2/3

Equilibrium price = 2/3